OpenStudy (anonymous):

I am having trouble with -4/5 + 7/15 = ?

OpenStudy (kainui):

The main thing to worry about with fractions is the denominator (the part on the bottoms). If you look at 15 and 5, what do they have in common? Well since 15 is just 5*3 it would be easier to combine them. For example, if I give you a half of a cake and a fourth of a cake, you can't just add 1/2+1/4 but if you just draw a picture you can see that a half of a cake with a fourth of a cake is just 3/4 of a cake, just look at this picture: So how do we make our numbers do this kind of thing? We want 1/2+1/4 to equal 3/4 since we see that it DOES! So what can we do? We can multiply anything by 1 and it doesn't change it. But if we're clever, we can change the denominator on the half. \[\Large \frac{1}{2}=\frac{1}{2}*1 =\frac{1}{2} * \frac{2}{2} = \frac{1*2}{2*2}=\frac{2}{4}\] See how 2/2=1? We didn't change the number, only how it appears. This lets us turn one half into two fourths so that we can see that two fourths plus one fourth equals 3 fourths! Try to do something similar here.

OpenStudy (anonymous):

but i am confused what to do about the negative fraction

OpenStudy (anonymous):

I Think the answer is 9/15 but i am still not sure

OpenStudy (kainui):

Show me your steps and I'll help you figure out where you went wrong. =)

OpenStudy (anonymous):

Wait never mind i messed up

OpenStudy (anonymous):

ok how about 1/3 that is my new answer

OpenStudy (anonymous):

-1/3

OpenStudy (kainui):

Yeah good job you got it

OpenStudy (anonymous):

Thank you!

OpenStudy (kainui):

You're welcome, keep it up =P